# A. Abed Abud CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland B. Abi

### v1

@@ VI Performance of reconstruction algorithm on supernova simulations
Supernova bursts are simulated to evaluate the performance of the reconstruction algorithm. The simulation is carried out by randomly picking eES and CC events from a large pool of events with uniformly distributed neutrino directions. The number of events selected per burst is based on the expected number of interactions for a 10 kpc supernova, as calculated by SNOwGLoBES (see Tab. 1). For each selected event, the reconstructed direction is rotated such that all events correspond to the same neutrino direction. To account for the effect of the anisotropic direction reconstruction performance, events are only selected if their neutrino direction lies within a cone of a 10 degree opening angle of the randomly selected supernova direction. In total, 10,000 supernova bursts are simulated and reconstructed using only eES events, and a pointing resolution (as defined in the previous section) of 3.4 degrees for a fiducial mass of 40 kton (four modules, full setup) and 6.6 degrees for one module of 10 kton is achieved.
The distribution of the truth-to-reconstruction angular difference for the burst is shown in Fig. 12 for two different assumptions on event classification quality. It is worth noting that only 0.1% of the simulated bursts’ reconstructed directions differed more than 90 degrees from the supernova position thanks to the brems flipping for both classification cases. The reconstructed directions of poorly reconstructed bursts are very close to being anti-parallel to the supernova direction.
Figure 12: Distribution of the angular difference between the reconstructed and true supernova direction, for 10,000 simulated supernova bursts. The distribution is shown for both perfect event classification and for an assumed 4% misclassification of CC events as eES as described in [37]. The ranges to the right of the respective colored dashed lines correspond to the 68% confidence intervals. The few bursts with a flipped reconstructed direction are excluded from the figure. The top figure corresponds to 10 kton fiducial mass, while it is 40 kton for the bottom figure.
@@
In adopting a more realistic classifier taking into account the misidentification of events, the aforementioned reconstruction procedure is effectively unchanged, except that the PDF used in (4) no longer corresponds to , but is rather a sum of and , weighted by the expected number of events from each interaction in the classification channel, which should be known ahead of time. As pointing information is primarily accessible from eES events, only the eES classification channel is used for direction reconstruction. Subsequently, the only elements in the confusion matrix that affect the pointing resolution are and . Figure 13 shows the pointing resolution as a function of these two matrix elements. For the worst-case scenario (where both elements are equal to ), the pointing resolution is degrees. While highly precise classifiers would provide highly accurate pointing results, even a weak classifier can yield a useful pointing resolution. At , only 20% better than random selection, the pointing resolution reaches around 40 degrees, significant enough to determine the quadrant of the sky that the supernova belongs to. An optimistic estimate of a boosted-decision-tree-based classifier is established in [37] at and . With this classifier, the pointing resolution amounts to degrees (8.7 degrees) for a fiducial mass of 40 kton (10 kton). The distribution of the reconstruction angle is compared to the perfect classification case in Fig. 12. While this study provides a rough estimation of the classification capabilities, further validation work is required to understand the classification performance in the presence of noise. Furthermore, because classification performance likely depends on the energy of the event, additional improvements in pointing resolution could be made by leveraging energy regions where the classifier performs best. Moreover, in a future fast online pipeline, real-time channel tagging before the comparison of the reconstructed events to the PDF will likely improve the overall performance as well.
Figure 13: Burst pointing resolution as a function of eES true positives () and CC false negatives (). Results assuming the fiducial mass of a single far detector module (10 kton) and all four planned modules (40 kton) are shown. For each pair of values, 1000 supernova bursts are simulated to determine the pointing resolution. Contour lines for various pointing resolution angles are also shown.
@@
Figure 14: Burst pointing resolution as a function of the number of detected eES events (), as well as the corresponding supernova distance. The event rates are calculated assuming the GKVM model at a given distance and are for a fiducial volume of 40 kton.
The supernova pointing resolution as a function of the supernova position in the sky is also examined. Fig. 15 shows the pointing resolution as a function of the detector coordinates. As for the results shown in Fig. 7, better pointing resolution can be seen closer to the directions (with being approximately the beam direction). Note that due to the way supernova bursts are simulated (as described at the beginning of this section), the dependence of the pointing resolution estimate is smeared around the incoming neutrino direction by the event selection radius of 10 degrees – but qualitatively, the variation in pointing resolution is of the order of a few degrees.
Fig. 16 shows the pointing resolution in the equatorial celestial coordinate system (RA/Dec), as a function of declination, averaged over right ascension. The figure also depicts the expected declination distribution for galactic supernovae to illustrate the most likely directions for a supernova to occur. Notably, because the direction of the detector is positioned at around of declination, the resolution is best around the same angle.
Figure 15: Burst pointing resolution as a function of the direction of the supernova, given in the detector coordinate system ( is the drift direction, is approximately the beam direction). (shown vertically) and (shown horizontally) are spherical coordinates in a coordinate system for which corresponds to the direction and correspond to the direction. Pointing resolution is given for the fiducial volumes of 10 kton and 40 kton. Perfect classification between eES and CC events is assumed in these plots. Note that the local pointing resolution is smeared by the event selection radius of 10 degrees.Figure 16: Burst pointing resolution as a function of the supernova direction, shown as a function of declination. The pointing resolution is averaged over right ascension. Shown on a separate axis is the expected declination distribution for galactic supernovae [38].
@@ VII Summary and outlook
An analysis framework for reconstructing supernova neutrino burst directions at DUNE is described. By reconstructing the eES and CC events during a supernova burst, the supernova direction can be determined from the correlation between the supernova neutrino and the outgoing primary electron created in the interaction with the LAr. The dominant interaction channels are CC and eES, with only the latter containing significant accessible information on the neutrino direction, but being an order of magnitude less frequent than the former.
Monte Carlo simulations in LArSoft were used to generate a data set of supernova bursts to evaluate the performance of the reconstruction algorithm for the planned DUNE detector setup.
From this data set, the energy and direction of the primary electron were reconstructed. To reduce head-tail ambiguities a highly effective technique called brems flipping was developed.
Finally, the pointing resolution is derived for ensembles of supernova burst events using a maximum likelihood method. With perfect event classification of the two event types considered, the pointing resolution is 3.4 degrees (6.6 degrees) with 68% coverage for supernovae at a distance of 10 kpc and an effective fiducial volume of 40 kton (10 kton). For a moderately optimistic classification performance, incorrect classification of 4% CC events as eES, the estimated pointing resolution is 4.3 degrees (8.7 degrees) with 68% coverage. The results presented here represent an average over all supernova directions in the sky.
In reality, there is a minor anisotropy in the burst reconstruction capability of the LAr TPC, resulting in a variation of the order of a few degrees in resolution.
@@
A continued effort should be carried out to further improve the current supernova pointing capabilities of DUNE. Prompt dissemination of directional information will be critical for multi-messenger astronomy, and efforts are underway to enable low-latency pointing with DUNE. Furthermore, given a core collapse at the most frequently expected distance from Earth for a galactic supernova, detection of supernova neutrino events can be expected in multiple large-scale neutrino detectors. By combining DUNE’s data with events in other detectors, one may take advantage of greater event statistics as well as the strengths of different detector technologies, resulting in a noticeable improvement to the current pointing resolution. Furthermore, we anticipate that steady improvements to pattern recognition technologies, especially via machine-learning techniques, will enhance the performance of next-generation particle track reconstruction and head-tail disambiguation algorithms, allowing for additional gains in pointing ability. Extended investigation of the eES/CC classification at DUNE is required, as this currently provides the greatest uncertainty in the pointing resolution. Furthermore, subdominant interaction channels should be considered as well. As a final note, laboratory investigation of neutrino interactions in the few-tens-of-MeV range in argon will be essential to fully understand the detector directional response [39].
@@ Acknowledgements
The ProtoDUNE-XX detector was constructed and operated on the CERN Neutrino Platform.
We gratefully acknowledge the support of the CERN management, and the
CERN EP, BE, TE, EN and IT Departments for NP04/ProtoDUNE-SP.
This document was prepared by the DUNE collaboration using the
resources of the Fermi National Accelerator Laboratory
(Fermilab), a U.S. Department of Energy, Office of Science,
HEP User Facility. Fermilab is managed by Fermi Research Alliance,
LLC (FRA), acting under Contract No. DE-AC02-07CH11359.
This work was supported by
CNPq,
FAPERJ,
FAPEG and
FAPESP, Brazil;
CFI,
IPP and
NSERC, Canada;
CERN;
MŠMT, Czech Republic;
ERDF,
H2020-EU and
MSCA, European Union;
CNRS/IN2P3 and
CEA, France;
INFN, Italy;
FCT, Portugal;
NRF, South Korea;
CAM,
Fundación “La Caixa”,
Junta de Andalucía-FEDER,
MICINN, and
Xunta de Galicia, Spain;
SERI and
SNSF, Switzerland;
TÜBİTAK, Turkey;
The Royal Society and
UKRI/STFC, United Kingdom;
DOE and
NSF, United States of America.
This research used resources of the
National Energy Research Scientific Computing Center (NERSC),
a U.S. Department of Energy Office of Science User Facility
operated under Contract No. DE-AC02-05CH11231.
@@ Appendix A Appendix
input : Two candidate vertices with positions , and momentum directions , ;

### v2

@@ VI Performance of reconstruction algorithm on supernova simulations
Supernova bursts are simulated to evaluate the performance of the reconstruction algorithm. The simulation is carried out by randomly picking eES and CC events from a large pool of events with uniformly distributed neutrino directions. The number of events selected per burst is based on the expected number of interactions for a 10 kpc supernova, as calculated by SNOwGLoBES (see Tab. 1). For each selected event, the reconstructed direction is rotated in such a way that the true burst directions align for all events. To account for the effect of the anisotropic direction reconstruction performance, events are only selected if their neutrino direction lies within a cone of a 10 degree opening angle of the randomly selected supernova direction. In total, 10,000 supernova bursts are simulated. Assuming no mis-identification of event types and using only the reconstrucred eES events, and a pointing resolution (as defined in the previous section) of 3.4 degrees for a fiducial mass of 40 kton (four modules, full setup) and 6.6 degrees for one module of 10 kton is achieved. The distribution of the truth-to-reconstruction angular difference for the burst is shown in Fig. 13 for two different assumptions on event classification quality. It is worth noting that only 0.1% of the simulated bursts’ reconstructed directions differed more than 90 degrees from the supernova position thanks to the brems flipping for both classification cases. The reconstructed directions of poorly reconstructed bursts are very close to being anti-parallel to the supernova direction.
@@
In adopting a more realistic classifier taking into account the misidentification of events, the aforementioned reconstruction procedure is effectively unchanged, except that the PDF used in (4) no longer corresponds to , but is rather a sum of and , weighted by the expected number of events from each interaction in the classification channel, which should be known ahead of time. As pointing information is primarily accessible from eES events, only the eES classification channel is used for direction reconstruction. Subsequently, the only elements in the confusion matrix that affect the pointing resolution are and . Figure 14 shows the pointing resolution as a function of these two matrix elements. For the worst-case scenario (where both elements are equal to ), the pointing resolution is degrees, averaging over a flat distribution of supernova directions. While highly precise classifiers would provide highly accurate pointing results, even a weak classifier can yield a useful pointing resolution. At and , only 20% better than random selection, the average pointing resolution reaches around 40 degrees, significant enough to determine the quadrant of the sky that the supernova belongs to. An optimistic estimate of a boosted-decision-tree-based classifier is described in [erin_conley_2020_4122909] at and . With this classifier, the average pointing resolution amounts to degrees (8.7 degrees) for a fiducial mass of 40 kton (10 kton). The distribution of the reconstruction angle is compared to the perfect classification case in Fig. 13. While this study provides a rough estimation of the classification capabilities, further validation work is required to understand the classification performance in the presence of noise. Furthermore, because classification performance likely depends on the energy of the event, additional improvements in pointing resolution could be made by leveraging energy regions where the classifier performs best. Moreover, in a future fast online pipeline, real-time channel tagging before the comparison of the reconstructed events to the PDF will likely improve the overall performance as well.
@@
The supernova pointing resolution as a function of the supernova position in the sky is also examined. Fig. 16 shows the pointing resolution as a function of the detector coordinates. As for the results shown in Fig. 8, better pointing resolution can be seen closer to the directions (with being approximately the beam direction). Note that due to the way supernova bursts are simulated (as described at the beginning of this section), the dependence of the pointing resolution estimate is smeared around the incoming neutrino direction by the event selection radius of 10 degrees – but qualitatively, the variation in pointing resolution is of the order of a few degrees. Fig. 17 shows the pointing resolution in the equatorial celestial coordinate system (RA/Dec), as a function of declination, averaged over right ascension. The figure also depicts the expected declination distribution for galactic supernovae to illustrate the most likely directions for a supernova to occur. Notably, because the direction of the detector is positioned at around of declination, the resolution is best around the same angle.
@@ VII Summary and outlook
An analysis framework for reconstructing supernova neutrino burst directions at DUNE is described. By reconstructing the eES and CC events during a supernova burst, the supernova direction can be determined from the correlation between the supernova neutrino and the outgoing primary electron created in the interaction with the LAr. The dominant interaction channels are CC and eES, with only the latter containing significant accessible information on the neutrino direction, but being an order of magnitude less frequent than the former. Monte Carlo simulations in LArSoft were used to generate a data set of supernova bursts to evaluate the performance of the reconstruction algorithm for the planned DUNE detector setup. From this data set, the energy and direction of the primary electron were reconstructed. To reduce head-tail ambiguities a highly effective technique called brems flipping was developed. Finally, the pointing resolution is derived for ensembles of supernova burst events using a maximum likelihood method. With perfect event classification of the two event types considered, the pointing resolution is 3.4 degrees (6.6 degrees) with 68% coverage for supernovae at a distance of 10 kpc and an effective fiducial volume of 40 kton (10 kton). For a moderately optimistic classification performance, incorrect classification of 4% CC events as eES, the estimated pointing resolution is 4.3 degrees (8.7 degrees) with 68% coverage. The results presented here represent an average over all supernova directions in the sky. In reality, there is a modest anisotropy in the burst reconstruction capability of the LAr TPC, resulting in a variation of the order of a few degrees in resolution.
@@
A continued effort will be carried out to further improve the current supernova pointing capabilities of DUNE. Prompt dissemination of directional information will be critical for multi-messenger astronomy, and efforts are underway to enable low-latency pointing with DUNE. Furthermore, given a core collapse at the most frequently expected distance from Earth for a galactic supernova, detection of supernova neutrino events can be expected in multiple large-scale neutrino detectors. By combining DUNE’s data with events in other detectors, one may take advantage of greater event statistics as well as the strengths of different detector technologies, resulting in a noticeable improvement to the current pointing resolution. Furthermore, we anticipate that steady improvements to pattern recognition technologies, especially via machine-learning techniques, will enhance the performance of next-generation particle track reconstruction and head-tail disambiguation algorithms, allowing for additional gains in pointing ability. Extended investigation of the eES/CC classification at DUNE is required, as this currently provides the greatest uncertainty in the pointing resolution. Furthermore, subdominant interaction channels should be considered as well. As a final note, laboratory investigation of neutrino interactions in the few-tens-of-MeV range in argon will be essential to fully understand the detector directional response [Asaadi:2022ojm].
@@ Acknowledgements
The ProtoDUNE-XX detector was constructed and operated on the CERN Neutrino Platform. We gratefully acknowledge the support of the CERN management, and the CERN EP, BE, TE, EN and IT Departments for NP04/Proto-DUNE-SP. This document was prepared by DUNE collaboration using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, Office of High Energy Physics HEP User Facility. Fermilab is managed by Fermi Forward Discovery Group, LLC, acting under Contract No. 89243024CSC000002. This work was supportedby CNPq, FAPERJ, FAPEG and FAPESP, Brazil; CFI, IPP and NSERC, Canada; CERN; MŠMT, Czech Republic; ERDF, Horizon Europe, MSCA and NextGenerationEU, European Union; CNRS/IN2P3 and CEA, France; INFN, Italy; FCT, Portugal; NRF, South Korea; Generalitat Valenciana, Junta de Andalucıa-FEDER, MICINN, and Xunta de Galicia, Spain; SERI and SNSF, Switzerland; TÜBİTAK, Turkey; The Royal Society and UKRI/STFC, United Kingdom; DOE and NSF, United States of America. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231.